OpenStudy (anonymous):

consider the equation hx+fy=c. Solve for x in terms h, f, c, and y. Then find the value of x if h=2, f=3, y=-5 and c=3

OpenStudy (anonymous):

hx=c-fy x=c-fy/h

OpenStudy (anonymous):

solving for x... x=3-(3)(-5)/2 =3+15/2 =18/2 =9